# CTR-KAN: KAN for Adaptive High-Order Feature Interaction Modeling

### v4

@@ Abstract
Modeling high-order feature interactions is critical for click-through rate (CTR) prediction, yet traditional approaches often face challenges in balancing predictive accuracy and computational efficiency. These methods typically rely on pre-defined interaction orders, which limit flexibility and require extensive prior knowledge. Moreover, explicitly modeling high-order interactions can lead to significant computational overhead. To tackle these challenges, we propose CTR-KAN, an adaptive framework for efficient high-order feature interaction modeling. CTR-KAN builds upon the Kolmogorov-Arnold Network (KAN) paradigm, addressing its limitations in CTR prediction tasks. Specifically, we introduce key enhancements, including a lightweight architecture that reduces the computational complexity of KAN and supports embedding-based feature representations. Additionally, CTR-KAN integrates guided symbolic regression to effectively capture multiplicative relationships, a known challenge in standard KAN implementations. Extensive experiments demonstrate that CTR-KAN achieves state-of-the-art predictive accuracy with significantly lower computational costs. Its sparse network structure also facilitates feature pruning and enhances global interpretability, making CTR-KAN a powerful tool for efficient inference in real-world CTR prediction scenarios. The source code is available at  [^1] https://github.com/Ancientshi/KarSein.

### v5

# Beyond KAN: Introducing KarSein for Adaptive High-Order Feature Interaction Modeling in CTR Prediction

@@ Abstract
Modeling high-order feature interactions is crucial for click-through rate (CTR) prediction, and traditional approaches often predefine a maximum interaction order and rely on exhaustive enumeration of feature combinations up to this predefined order. This framework heavily relies on prior domain knowledge to define interaction scope and entails high computational costs from enumeration. Conventional CTR models face a trade-off between improving representation through complex high-order feature interactions and reducing computational inefficiencies associated with these processes. To address this dual challenge, this study introduces the Kolmogorov-Arnold Represented Sparse Efficient Interaction Network (KarSein). Drawing inspiration from the learnable activation mechanism in the Kolmogorov-Arnold Network (KAN), KarSein leverages this mechanism to adaptively transform low-order basic features into high-order feature interactions, offering a novel approach to feature interaction modeling. KarSein extends the capabilities of KAN by introducing a more efficient architecture that significantly reduces computational costs while accommodating two-dimensional embedding vectors as feature inputs. Furthermore, it overcomes the limitation of KAN’s its inability to spontaneously capture multiplicative relationships among features. Extensive experiments highlight the superiority of KarSein, demonstrating its ability to surpass not only the vanilla implementation of KAN in CTR prediction tasks but also other baseline methods. Remarkably, KarSein achieves exceptional predictive accuracy while maintaining a highly compact parameter size and minimal computational overhead. Moreover, KarSein retains the key advantages of KAN, such as strong interpretability and structural sparsity. As the first attempt to apply KAN in the CTR domain, this work introduces KarSein as a novel solution for modeling complex feature interactions, underscoring its transformative potential in advancing CTR prediction task. The code is available at  [^1] https://github.com/Ancientshi/KarSein.